OpenStudy (anonymous):

y=x-2 and x=-y-4 using substitution method.

OpenStudy (solomonzelman):

OK! you got y=x-2 and x=-y-4 you know what x is in terms of y, you are given that strait out x=-y-4 so plug -y-4 for x.

OpenStudy (solomonzelman):

y=x-2 y=(-y-4)-2 y=-y-4-2 y=-y-6 2y=-6 y=-3

OpenStudy (anonymous):

its -1

OpenStudy (solomonzelman):

what is the x? yes!

OpenStudy (solomonzelman):

did this help you?

OpenStudy (anonymous):

at this part y=-y-6 2y=-6 y=-3 how is it 2y??

OpenStudy (solomonzelman):

b/c you add y to both sides, (or bring -y to the other side) to eliminate it.

OpenStudy (solomonzelman):

Makes sense, right?

OpenStudy (anonymous):

yep! :D

OpenStudy (anonymous):

how to graph it?

OpenStudy (solomonzelman):

the same problem? put it in a y=mx+b form and plug in any numbers for x.

OpenStudy (solomonzelman):

(I don't recommend plugging in really big numbers, try 1,2 and connect the points with a line)

OpenStudy (anonymous):

Thank You, Solomon! I was unsure how to use the substitution on this, I know how to graph it.